# VisualLens: Personalization through Visual History

### v1

@@ 5.1 Multi-image caption pretraining
To facilitate the model to ground to each grid faithfully, we perform a LoRA continual pretraining on dense captions. We adopt the dense captioning dataset DOCCI dataset (Onoe et al., 2024), which contains over 15,000 images and their corresponding dense captions. Each time, we randomly sample $w$ images =$\mathcal{I}=\{i_{1},\cdots,i_{w}\}$ and their corresponding caption $\mathcal{C}=\{x_{1},\cdots,x_{w}\}$, and then we construct a gridified input image $G(\mathcal{I})$ and a target output text description $T(\mathcal{C})=$Image 1: $x_{1}$, $\cdots$, Image w: $x_{w}$”. Then we LoRA finetune the pretrained backbone model (e.g., MiniCPM-V2.5) on all image-caption pairs $\{G(\mathcal{I}),T(\mathcal{C})\}$ so that the model is able to process the gridified user history grid by grid. We then use the continual pretrained model as the starting point to apply joint training described in §5.3.
5.2 Iterative aspect word refinement
@@
Unlike image captioning, aspect word generation is not a standard multimodal task and lacks the extensive pretraining data typically available. Hence, with zero-shot prompting, the generated aspect words have a large quality gap across different images, and the extracted aspects may not indicate user preferences.
To finetune aspect word generation, we first generate the training data. For each image $i$, we start from an initial set of aspect words, denoted as $\mathcal{W}_{i}^{(0)}$, which is generated by LLaVA-v1.6. In the $j^{\text{th}}$ round, we prompt a separate Llama-3.1 70B model with $\mathcal{W}_{i}^{(j-1)}$, candidates, and ground truths, and ask it to select useful aspect words that are helpful in ground truth prediction, which constitute $\mathcal{W}_{i}^{(j)}$. This refinement process continues for several rounds, and the iterations allow for converging extracted aspect words toward a more accurate and relevant subset. Empirically, we observe that the refinement converges after approximately 4 rounds, and denote ${}^{\textnormal{th}}$ refined aspect word set $\mathcal{W}_{i}^{(4)}$ as $\mathcal{W}_{i}$, which serves as the training target.
@@
5.3 Joint training of aspect word generation and candidate matching
@@
where $u_{j},q_{j},\mathcal{S}_{j}$ is the user, the question, and the ground truth set of candidates of the $j$-th example. The text prompt $x_{\text{pred},j}$ consists of the question $q_{j}$, the text user preference profiles $x_{u_{j},q_{j}}$, and candidates. We LoRA finetune the model under the joint loss $\mathcal{L}_{\text{joint}}$.
{NiceTabular}
[…]
GR-V 15.69M 2K 200K 66 157.0 2.7 43.1
Yelp-V 4.12M 2K 100K 35 263.6 8.2 66.7
Table 1: Dataset statistics of Google Review-Vision (GR-V) and Yelp-Vision (Yelp-V).
@@
Table 2: Hit rates and MRRs of VisualLens vs. multiple baselines on Google Review-V and Yelp-V. The result shows (a) VisualLens outperforms other baselines, though has a gap with the human oracle; (b) model size greatly affects the performance; (c) simply rank by rating is a worse design than the random baseline.
@@ 6.1 Benchmark creation
To the best of our knowledge, there is no existing benchmark (Harper and Konstan, 19; Ni et al., 2019; Wu et al., 2020; Wan and McAuley, 2018; Salemi et al., 2023) to evaluate personalization with task-agnostic visual history. We created two benchmarks, Google Review-V and Yelp-V, leveraging publicly available data from Google Local Review (Li et al., 2022) and Yelp (Asghar, 2016).
User logs:
For each user in the two datasets, we take the list of reviews in chronological order. Each review is associated with a business name, categories and description. In Google Review-V, each review is associated with a few photos, used for image logs. Yelp-V does not associate a review with photos, so we randomly subsample one-third of the store profile pictures, such that different reviews for the same business can be associated with different images.
Questions and visual history:
We consider a special type of questions, category recommendations, like "Recommend a nearby museum". Such questions are both popular in real applications, and hard as there are many candidates satisfying the constraint. We remove small categories and most ambiguous categories, such as “place”, and “spot”.
For each review $r$ regarding a business of category $c$, we create a question asking to recommend businesses in category $c$. We take all (and only) photos in the reviews before $r$ to generate the visual history. We consider the visual history task-agnostic since the categories are highly diverse (see Figure 4, and the photos are quite diverse too (e.g., a park photo to illustrate happiness mentioned in the review). We filter instances where the history is too short ($<$$10$) or does not contain the questioned category.
Candidates and ground truths:
For a review $r$ we take all reviews starting from $r$ to generate candidates and ground truths. To be realistic, we consider only nearby businesses of the same category as the candidate set, and the number of candidates is a random number in $[30,100]$. Candidates that also appear in the user’s future reviews are considered as ground truths. To avoid falling into a classification problem, we filter examples with only 1 ground truth in Google Review-V and fewer than 5 in Yelp-V.
Summary:
Table 5.3 gives the benchmark statistics. The ratio of an average number of candidates and that of ground truths is 16:1 for Google Review-V and 8:1 for Yelp-V, showing the difficulties of recommendation. By default, the train, dev, and test data have disjoint users. We discuss other splitting in Table 6.3.
@@ 6.2 Evaluation measures
Mean Reciprocal Rank (MRR).
$\text{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\text{rank}(r_{i})}$ measures the ranking quality by averaging the reciprocal ranks of the first relevant item for each example. The MRR ranges from $1/S$ to $1$, where $S$ is the number of candidates.
@@ 6.3 Implementation and baselines
We ran VisualLens with two backbone models, a smaller 3B model PaliGemma and a larger 8B model MiniCPM-V2.5. We used an image grid of $8\times 8$, and a loss weighting factor $\lambda=2$ in Eqn. 3. We compared VisualLens with multiple other solutions.
@@
  • Fine-tuned models: We fine-tuned three state-of-the-art solutions: multimodal personalization model UniMP (Wei et al., 2024a) (RedPajama 3B model), with the adaptation to replace their item image and attributes with image tokens in the visual history; Llama-3.1-8B-Instruct (Meta, 2024) with text-only user preference profiles; and MiniCPM-V2.5 8B (Yao et al., 2024a) with image profiles.
@@
Table 3: Ablation study on PaliGemma. Different components of VisualLens model: joint training (Joint), iterative refinement (Iter), aspect words (Asp.), captions (Cap.), image embedding (Img.), and relevant image retrieval (Ret.). MRR difference over 0.4 has a p-value$<$0.04.
@@
7 Results and Analysis
We conducted experiments to answer four questions:
  • Q3: Can VisualLens transfer across users, unseen categories and longer history?
  • Q4: What is the robustness of VisualLens?
7.1 Recommendation effectiveness (Q1)
VisualLens significantly outperforms baselines.
Our first observation from Table 5.3 is that all recommendation models that leverage the visual history significantly outperform baseline solutions on all metrics. In particular, the best version of VisualLens improves Hit@3 over Random by 18% on Google Review-V and by 26% on Yelp-V, and even more over Rank by rating (apparently, overall ratings do not cater to specific users). There is still a gap between VisualLens and human annotations, but comparing w. Random, it fills $\sim$75of the gaps for hit@3 on both datasets.
VisualLens outperforms state-of-the-art solutions.
VisualLens outperforms UniMP with the same number of trainable parameters. With the 8B MiniCPM-V2.5 backbone, VisualLens outperforms MiniCPM-V2.5 itself by 2.5% on Hit@3 on Google Review-V, and by 6% on Yelp-V, and we observe a similar trend for the 3B models. Even comparing with significantly larger models, including Llama-3.1-70B-Instruct and GPT-4o without fine-tuning, VisualLens 7B improves by 1.6-5.6% on Hit@3.
7.2 Ablation studies (Q2)
We evaluate the usefulness of each component in VisualLens in Table 6.3, with PaliGemma as the backbone.
  1. 1.
    History retrieval significantly improves the results, and can reduce Hit@3 by 7% on Google Review-V and by 12% on Yelp-V (#7 vs. #3).
  2. 2.
    All three representations of the images (embedding, caption, aspects) play an important role. Removing captions and aspect words can reduce Hit@3 by 3% on Google Review-V and by 9% on Yelp-V, even without fine-tuning (#6 vs. #3). Between the two, aspect words play a more important role than captions (#5 vs. #4).
  3. 3.
    Both iterative training and joint multi-task training improve the recommendation quality. Removing both of them lowers Hit@3 by 1%+ on both data sets (#3 vs #1).
@@ 7.3 Transferability (Q3)
We tested transferability over users (default setting), over categories, and over different (longer) history lengths. Table 6.3 compares the MRR of VisualLens on both benchmarks, showing good transferability. MRR is highest when applied to longer history, with the effectiveness of history retrieval. Transferability is higher across categories than across users, both demonstrating the promise of leveraging task-agnostic signals, and illustrating slightly more challenges to transfer between users of different interest patterns.
7.4 Robustness (Q4)
Finally, we conducted a few robustness tests.
  1. 1.
    Figure 3(a) shows that as the number of candidates grows, the recommendation becomes harder and MRR gets lower. However, when the number of candidates exceeds 50, MRR converges.
  2. 2.
    Figure 3(b) shows that as the history grows, MRR increases and flattens after reaching $\sim$100 images. This trend is related to our grid size $8\times 8$, since smaller than 64 images would not leverage all spaces in the grid. On the other hand, flat MRR after 100 images shows robustness against history noises with the retrieval step.
  3. 3.
    Finally, Figure 4 plots MRR for different categories. There are a few factors that can affect recommendation quality. First and foremost, ambiguous categories like “area”, “station”, and “market” get the lowest MRR in Google Review-V. Second, general categories (e.g., “museum”, “hotel”) with bigger category sizes obtain higher MRR than specific ones (e.g., “historical landmark”). Third, transfer learning happens between neighbor categories; for example, “deli” and “takeout” achieve the top-$2$ and top-$3$ performance with less training data, since they are similar to the largest category “restaurant”.
@@ 8 Conclusion and Discussion
In this paper, we proposed a novel approach VisualLens for personalized recommendation using task-agnostic visual history. We advanced multimodal LLMs to extract meaningful signals from images to serve as proxies of user preferences. We created two benchmarks, Google Review-V and Yelp-V, to evaluate VisualLens, affirming the efficacy and robustness of VisualLens.
VisualLens lends itself as a first step towards recommendation systems leveraging task-agnostic visual history. There are still many future directions to explore. First, we can integrate the VisualLens solution with other accessible information, such as the date and location each image is taken, the specific fine-granularity entities (e.g., particular products) recognized from the images, and other user profiles. Second, we can extend the recommendation problems studied in this paper with broader QA tasks. Finally, we shall study how to guarantee privacy in the training with techniques like federated learning.
@@ A.1 Algorithm for candidate generation
Input: Business geographics graph $\mathcal{G}=(\cal{V},\cal{E})$, Visit set $\cal{B}$ of a user, Category $c$, Minimum random candidate count ${rc}_{\text{min}}$, Maximum random candidate count ${rc}_{\text{max}}$, Minimum final candidate count ${fc}_{\text{min}}$, Minimum ground truth count ${gtc}_{\text{min}}$.
@@ A.2 Processing and filtering
Category removal and merge.
@@ B More Related Works
B.2 Datasets and evaluating multimodal recommendation.
@@
C Prompt Template

### v2

@@ 4.3.1 Multi-image caption pretraining
To facilitate the model to ground to each grid faithfully, we perform a LoRA continual pretraining on dense captions. We adopt the dense captioning dataset DOCCI dataset (Onoe et al., 2024), which contains over 15,000 images and their corresponding dense captions. Each time, we randomly sample $w$ images $\mathcal{I}=\{i_{1},\cdots,i_{w}\}$ and their corresponding caption $\mathcal{C}=\{x_{1},\cdots,x_{w}\}$, and then we construct a gridified input image $G(\mathcal{I})$ and a target output text description $T(\mathcal{C})=$Image 1: $x_{1}$, $\cdots$, Image w: $x_{w}$”. Then we LoRA finetune the pretrained backbone model (e.g., MiniCPM-V2.5) on all image-caption pairs $\{G(\mathcal{I}),T(\mathcal{C})\}$ so that the model is able to process the gridified user history grid by grid. We then use the continual pretrained model as the starting point to apply joint training described in §4.3.3.
4.3.2 Iterative aspect word refinement
@@
Unlike image captioning, aspect word generation is not a standard multimodal task and lacks the extensive pretraining data. Hence, with zero-shot prompting, the generated aspect words have a large quality gap across different images, and the extracted aspects may not indicate user preferences.
[…]
@@
4.3.3 Joint training of aspect word generation and candidate matching
@@
where $u_{j},q_{j},\mathcal{S}_{j}$ is the user, the question, and the ground truth set of candidates of the $j$-th example. The text prompt $x_{\text{pred},j}$ consists of the question $q_{j}$, the text query-specific profiles $x_{u_{j},q_{j}}$, and candidates. We LoRA finetune the model under the joint loss $\mathcal{L}_{\text{joint}}$.
Table 3: Hit rates and MRRs of VisualLens vs. multiple baselines on Google Review-V and Yelp-V. The result shows (a) VisualLens outperforms other baselines, though has a gap with the human oracle; (b) model size greatly affects the performance; (c) simply rank by rating is a worse design than the random baseline. Due to the large test set size (200K), an MRR difference greater than 0.4 yields a $p$-value << 0.04.
@@ 5.1 Benchmark creation
To the best of our knowledge, there is no existing benchmark (Harper and Konstan, 19; Ni et al., 2019; Wu et al., 2020; Wan and McAuley, 2018; Salemi et al., 2023) to evaluate personalization with task-agnostic visual history. We created two benchmarks, Google Review-V and Yelp-V, leveraging publicly available data from Google Local Review (Li et al., 2022) and Yelp (Asghar, 2016). Table 2 gives the benchmark statistics. The ratio of an average number of candidates and that of ground truths is 16:1 for Google Review-V and 8:1 for Yelp-V. By default, the train, dev, and test data have disjoint users. We discuss other splitting in Table 6.2. We list the details on dataset creation in Appendix A.
@@ 5.2 Evaluation measures
Mean Reciprocal Rank (MRR). $\text{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\text{rank}(r_{i})}$ measures the ranking quality by averaging the reciprocal ranks of the first relevant item for each example. The MRR ranges from $1/S$ to $1$, where $S$ is the number of candidates.
@@ 5.3 Implementation and baselines
We ran VisualLens with two backbone models, a smaller 3B model PaliGemma and a larger 8B model MiniCPM-V2.5. For optimal performance, we selected $w=64$ (112x112 each sub-image for PaliGemma, best in practice), corresponding to an $8\times 8$ image grid, a candidate count of n=10n=10k, and a loss weighting factor of $\lambda=2$ (Eqn. 3). We compared VisualLens with solutions below.
  • $\bullet$
@@
  • $\bullet$
    Fine-tuned models: We fine-tuned three state-of-the-art solutions: multimodal personalization model UniMP (Wei et al., 2024a) (RedPajama 3B), with the adaptation to replace item images and attributes with image tokens in the visual history; Llama-3.1-8B-Instruct (Meta, 2024) with text-only user preference profiles; PaliGemma 3B and MiniCPM-V2.5 8B (Yao et al., 2024) with image profiles.
  • $\bullet$
@@
Table 4: Ablation study on PaliGemma. Different components of VisualLens model: joint training (Joint), iterative refinement (Iter), aspect words (Asp.), captions (Cap.), image embedding (Img.), and relevant image retrieval (Ret.). An MRR difference greater than 0.4 yields a $p$-value << 0.04.
@@
6 Results and Analysis
We conducted experiments to answer four questions:
  • Q3: Can VisualLens transfer across users, unseen categories, longer history, and new benchmark?
  • Q4: What is the robustness of VisualLens?
6.1 Recommendation effectiveness (Q1)
VisualLens significantly outperforms baselines. Our first observation from Table 4.3.3 is that all recommendation models that leverage the visual history significantly outperform baseline solutions on all metrics. In particular, the best version of VisualLens improves Hit@3 over Random by 18% on Google Review-V and by 26% on Yelp-V, and even more over Rank by rating (apparently, overall ratings do not cater to specific users). There is still a gap between VisualLens and human annotations, but comparing w. Random, it fills $\sim$75% of the gaps for hit@3 on both datasets.
VisualLens outperforms state-of-the-art solutions. VisualLens outperforms UniMP with the same number of trainable parameters. With the 8B MiniCPM-V2.5 backbone, VisualLens outperforms MiniCPM-V2.5 itself by 2.5% on Hit@3 on Google Review-V, and by 6% on Yelp-V, and we observe a similar trend for the 3B models. Even comparing with significantly larger models, including Llama-3.1-70B-Instruct and GPT-4o without fine-tuning, VisualLens 7B improves by 1.6-5.6% on Hit@3.
6.2 Ablation studies (Q2)
We evaluate the usefulness of each component in VisualLens in Table 5.3, with PaliGemma as the backbone. We find history retrieval significantly improves the results, and can reduce Hit@3 by 7% on Google Review-V and by 12% on Yelp-V (#7 vs. #3). Besides, all three representations of the images (embedding, caption, aspects) play an important role. Removing captions and aspect words can reduce Hit@3 by 3% on Google Review-V and by 9% on Yelp-V, even without fine-tuning (#6 vs. #3). Between the two, aspect words play a more important role than captions (#5 vs. #4). Moreover, both iterative training and joint multi-task training improve the recommendation quality. Removing both of them lowers Hit@3 by 1%+ on both data sets (#3 vs #1).
@@ 6.3 Transferability (Q3)
We tested transferability over users (default setting), over categories, and over different (longer) history lengths. Table 6.2 compares the MRR of VisualLens on both benchmarks, showing good transferability. MRR is highest when applied to longer history, with the effectiveness of history retrieval. Transferability is higher across categories than across users, both demonstrating the promise of leveraging task-agnostic signals, and illustrating slightly more challenges to transfer between users of different interest patterns. We also present the generalization results across benchmarks in Appendix C.
6.4 Robustness and qualitative analysis (Q4)
We conducted several robustness tests as follows.
Candidate count: Figure 2(a) shows that as the number of candidates grows, the recommendation becomes harder and MRR gets lower. However, when the number of candidates exceeds 50, MRR converges.
Image count in user history: Figure 2(b) shows that as the history grows, MRR increases and flattens after reaching $\sim$100 images. This trend is related to our grid size $8\times 8$, as smaller than 64 images will not leverage all spaces in the grid. On the other hand, flat MRR after 100 images shows robustness against history noises with the retrieval step.
Category distribution: Figure 4 plots MRR for different categories. There are a few factors that can affect recommendation quality. First and foremost, ambiguous categories like “area”, “station”, and “market” get the lowest MRR in Google Review-V. Second, general categories (e.g., “museum”, “hotel”) with bigger category sizes obtain higher MRR than specific ones (e.g., “historical landmark”). Third, transfer learning happens between neighbor categories; for example, “deli” and “takeout” achieve the top-$2$ and top-$3$ performance with less training data, since they are similar to the largest category “restaurant”. We provide more qualitative analysis in Appendix C.
@@ 7 Conclusion and Discussion
In this paper, we proposed a novel approach VisualLens for personalized recommendation using task-agnostic visual history. We advanced MLLMs to extract spectrum signals from images to serve as user profile. We created two benchmarks, Google Review-V and Yelp-V, to evaluate VisualLens, affirming the efficacy and robustness of VisualLens.
VisualLens offers a promising first step toward MLLM recommendation systems that leverage task-agnostic visual history. Several future directions remain. First, we could integrate VisualLens with additional data, such as image timestamps, locations, recognized fine-grained entities (e.g., specific products), and user profiles. Second, we aim to extend the recommendation problems explored here to encompass broader QA tasks. Finally, we plan to investigate privacy-preserving techniques, such as federated learning, during training.
Limitation
While VisualLens demonstrates strong performance on task-agnostic multimodal recommendation, several limitations remain.
Modular design without optimal components.
Our framework prioritizes modularity and extensibility, using a unified architecture across modules including image encoding, captioning, aspect word generation, and preference matching. However, we do not use the best-performing model for each subtask. Each component can be replaced with more advanced or domain-specific alternatives, which could potentially boost overall performance.
Limited modality and domain coverage.
Our experiments focus exclusively on static visual history (images) and do not cover more complex modalities such as video, audio, or multimodal narratives over time. These richer formats are important for building truly comprehensive, lifelong user models, as envisioned by early ideas like Memex (Bush, 1945), but remain out of scope for this work.
Evaluation scope.
We restrict the recommendation task to a QA format for consistency and interpretability. While this setup allows for controlled benchmarking and comparison, it does not cover other common recommendation forms such as ranked lists, interactive dialogues, or implicit feedback modeling. Additionally, our proposed benchmarks, Google Review-V and Yelp-V, capture a subset of the full task-agnostic recommendation landscape and may not reflect the full range of real-world personalization scenarios.
Future work could explore richer modalities, more diverse recommendation formats, and further improvements to the modular components of VisualLens.
@@ Acknowledgements
Social Impact
The ability to model user preferences from visual histories raises important considerations around privacy, consent, and data usage. While our work uses publicly available and anonymized images, real-world deployments would need to carefully address how user data is collected, stored, and interpreted, especially when recommendations extend beyond a single domain. There is also potential for reinforcing biases or overfitting to superficial cues (e.g., location, aesthetics) that may not reflect deeper user intent. We encourage future research to incorporate fairness auditing, privacy-preserving training, and transparency mechanisms to ensure responsible use of multimodal personalization systems.
@@ A.1 Algorithm for candidate generation
[…]
@@ A.2 Processing and filtering
User logs: For each user in the two datasets, we take the list of reviews in chronological order. Each review is associated with a business name, categories and description. In Google Review-V, each review is associated with a few photos, used for image logs. Yelp-V does not associate a review with photos, so we randomly subsample one-third of the store profile pictures, such that different reviews for the same business can be associated with different images.
Category removal and merge:
@@
Questions and visual history: We consider a special type of questions, category recommendations, like "Recommend a nearby museum". Such questions are both popular in real applications, and hard as there are many candidates satisfying the constraint. We remove small categories and most ambiguous categories, such as “place”, and “spot”.
For each review $r$ regarding a business of category $c$, we create a question asking to recommend businesses in category $c$. We take all (and only) photos in the reviews before $r$ to generate the visual history. We consider the visual history task-agnostic since the categories are highly diverse (see Figure 4), and the photos are quite diverse too (e.g., a park photo to illustrate happiness mentioned in the review). We filter cases where the history is too short (<<1010) or does not contain the questioned category. By doing so, we ensure that the user history contains at least 10 relevant images (i.e., from the same category as the query) for each example in both Google Review-V and Yelp-V.
Candidates and ground truths: For a review $r$ we take all reviews starting from $r$ to generate candidates and ground truths. To be realistic, we consider only nearby businesses of the same category as the candidate set, and the number of candidates is a random number in $[30,100]$. Candidates that also appear in the user’s future reviews are considered as ground truths. To avoid falling into a classification problem, we filter examples with only 1 ground truth in Google Review-V and fewer than 5 in Yelp-V.
@@ B More Related Works
B.2 Multimodal large language models.
Multimodal LLMs are becoming increasingly powerful in processing images and videos and in generating human-like text, exemplified by models like GPT-4 family (OpenAI, 2024), Claude 3.5 family (Anthropic, 2024), Gemini and PaliGemma (Google, 2023; Beyer et al., 2024), LLaVA model family (Liu et al., 2024), and Llama 3 Vision models (Meta, 2024). However, they still suffer from a strong language prior (Fu et al., 2024; Parcalabescu and Frank, 2024), generalization (Zhu et al., 2022) and hallucinations (Fu et al., 2025), or require heavy text transcription (Zhu et al., 2024a). The extent to which they can understand user history, in analogy to LLM recommendation systems, is still unclear.
B.3 Datasets and evaluating multimodal recommendation.
@@
C More Analysis
C.1 Sensitivity of random sampling for candidate embedding
We show that the recommendation results are not sensitive to the random sampling of images in candidate embedding. Table C.1 below shows that the robustness against sampling randomness: 1) the standard deviation is <<0.1 when we do 5 runs of sampling; 2) as we reduce the size of the samples to 1K and 100, the MRR stays similar, with <<0.5 difference.
{NiceTabular}
[…]
10k 33.5 $\pm$ 0.1 44.3 $\pm$ 0.0
1k 33.4 $\pm$ 0.1 44.3 $\pm$ 0.1
100 33.0 $\pm$ 0.2 44.1 $\pm$ 0.1
Table 6: MRR and its standard deviation (5 runs) of VisualLens (PaliGemma) on different sampling sizes.
C.2 Generalization between Google Review-V and Yelp-V
We report the generalization quality between Google Review-V and Yelp-V. Table C.2 below shows that cross-source recommendation reduces MRR by 2-4% respectively vs. in-domain, in an acceptable range, highlighting the transferability from the shared feature in the embedding space. Besides, transferring from Google Review-V to Yelp-V is still better than the best baseline model in-domain (40.2), while the other side is much worse (33.2). This is as expected, since Google Review-V has much more categories and much lower photo quality.
{NiceTabular}
[…]
Google Review-V Yelp-V
Google Review-V 33.5 41.9
Yelp-V 29.4 44.3
Table 7: MRR of VisualLens (PaliGemma) on in-domain test and cross-domain transfer.
C.3 Qualitative results
We show the input and output of a positive example in Figure 6 and Figure 7. We show the input and output of a negative example in Figure 8 and Figure 9. Please note that due to image licensing restrictions, we are only able to provide the image names in JPG format. For more details, you can download the image data from https://business.yelp.com/data/resources/open-dataset/.
Figure 6: A successful example of VisualLens on Hit@1. We list the input question, images and candidates.
Figure 7: A successful example of VisualLens on Hit@1. We list the output candidate ranking and ground truth. The ranking follows the same left-right order as the input candidates.
Figure 8: A failed example of VisualLens on Hit@3. We list the input question, images and candidates.
Figure 9: A failed example of VisualLens on Hit@3. We list the output candidate ranking and ground truth. The ranking follows the same left-right order as the input candidates.
D Prompt Template